# Interpretable Graph Learning Over Sets of Temporally-Sparse Data

### v2

@@ 5 Conclusion
In this work, we present GMAN, a novel method that extends GNAN[4] for learning from sets of temporally sparse graphs, while allowing a flexible balance between interpretability and expressivity. We show that GMAN achieves state-of-the-art results across real-world medical predictions tasks with high-stake implications, as well as fake news detection, including over $4$ AUCROC points improvement in predicting in-hospital mortality. We demonstrate how the strong interpretability capabilities of GMAN, including node-, graph-, and set-level explanations, can be leveraged to uncover novel clinical insights, such as identifying transition phases in patient disease progression. We provide a theoretical analysis of GMAN, and show it is strictly more expressive than GNAN.
Overall, GMAN’s flexibility and expressiveness makes it well-suited for a wide range of real-world tasks involving irregular and heterogeneous data, particularly in domains such as healthcare where interpretability is crucial.
Appendix
6 Theoretical Framework
This section consolidates the theoretical foundations of GMAN by presenting formal proofs of its expressive power and structural recoverability guarantees. We first prove that GMAN is strictly more expressive than GNAN, both in terms of multivariate feature learning and subset level representation. We then establish conditions under which GMAN can recover the latent structure of input graphs purely from learned pairwise distances. Specifically, we show that when graphs form tree-structured trajectories, the model’s learned distance matrix implicitly determines the original graph topology, up to isomorphism.
6.1 Proof of Theorem 3.1
We will prove that GMAN is strictly more expressive than GNAN. To prove this, we use a ground truth function that is a feature-level XOR. Let a single-node graph be endowed with binary features $x=(x_{1},x_{2})\in\{0,1\}^{2}$ and define the target $f_{\oplus}(x)=x_{1}\oplus x_{2}$.
@@ 6 Theoretical Framework
Now we will show that GMAN can express $f_{\oplus}$. Place the two features in the same subset $F=\{x_{1},x_{2}\}$ and choose the subset-network
@@
For the four binary inputs this mapping returns $(0,1,1,0)$, exactly $f_{\oplus}$. Hence GMAN represents a function unattainable by GNAN, proving that GMAN is strictly more expressive.
6.2 Proof of Theorem 3.2
Let $S$ be a set of graphs $\{G_{i}\}_{j=1}^{m}$. Let $S_{1}=\{S_{i}\}_{i=1}^{m}$ be a partition of $S$ such that $|S_{i}|=1$. Let $S_{2}=\{S_{i}\}_{i=1}^{k}$ such that there exists $k$ with $|S_{k}|>1$. with a subset partition $\{S_{i}\}_{i=1}^{k}$. We will prove that a GMAN trained over $S_{2}$ is strictly more expressive than a GMAN trained over $S_{1}$.
@@
If each graph is placed in its own subset, GMAN aggregates additively: the model output is
@@
because the final GMAN stage simply sums subset scores :contentReference[oaicite:0]index=0:contentReference[oaicite:1]index=1. Write $a=\phi(0)$ and $b=\phi(1)$. To realise $f_{\oplus}$ via a threshold $\tau$ we would need
@@
Group the two graphs together and use a DeepSet $\Phi(S_{2})=g\!\bigl{(}\sum_{i=1}^{2}f(x_{i})\bigr{)}$ with $f(x)=x$ and $g(s)=s(2-s)$. Then
@@
6.3 Four-Point Condition and Recoverability
We now turn to structural identifiability. We prove that when input graphs are connected, acyclic, and positively weighted (i.e., trees), the pairwise distance matrix learned by GMAN encodes the full structure of the graph, up to isomorphism. This provides theoretical justification for the model’s ability to reason over temporal structure without needing explicit graph supervision.
The following theorem shows that if the graph satisfies the four-point condition [6], GMAN can reconstruct the original graph from the transformed distance matrix that is fed to GMAN as the graph input:
Theorem 6.1.
Let $G$ be a graph represented by an adjacency matrix $A$, and $D$ be the transformed distance-matrix for GMAN. Then if $D$ satisfies the four-point condition, GMAN can learn $\rho$ such that $\rho(D)=A$.
@@ 6.3 Four-Point Condition and Recoverability
Define the pair-wise distance matrix $D\in\mathbb{R}^{n\times n}$ by
@@
where $P_{G}(u,v)$ is the unique $u$$v$ path in $G$. Then:
@@
    Tree-metric property. $D$ satisfies the four-point condition of Buneman [6]; hence $(V,D)$ is a tree metric.
@@
7 Dataset Details
7.1 PhysioNet P12
@@
7.2 Crohn’s Disease Prediction
@@
7.2.1 Clinical Context and Related Work for predicting CD onset
@@
Several studies have assessed the predictive potential of standard clinical blood tests. For example, [45] analyzed six routine biomarkers from 1,186 Danish patients eventually diagnosed with CD, achieving moderate predictive performance (AUROC of 0.74) approximately six months before clinical diagnosis. Larger-scale analyses, such as those using UK Biobank data, combined multiple standard biomarkers and basic demographic information, reporting similar predictive performances (AUROCs typically between 0.70–0.75). These analyses generally utilized methods like logistic regression, random forests, or gradient-boosted trees, favored for structured clinical datasets.
Other studies have integrated advanced biochemical data, known as multi-omics, including large-scale protein measurements (proteomics), metabolites (metabolomics), or genomic markers. [15] for instance, combined 67 blood biomarkers with approximately 2,900 plasma proteins from the UK Biobank, achieving an AUROC of 0.786. [46] combined genetic risk scores with extensive proteomic data, achieving an AUROC of 0.76 for CD prediction up to five years prior to diagnosis. Similar multi-omics approaches employing microbiome profiling, immune signaling molecules (cytokines), or lipid molecules typically achieve AUROCs between 0.75 and 0.80 but often involve significant cost, specialized laboratory analyses, and reduced consistency across diverse patient cohorts.
@@
8 Biomarker Subset Groupings
8.1 P12
@@
    Respiratory gas exchange and ventilation
    [FiO2, PaO2, PaCO2, SaO2, RespRate, pH, MechVent]
    These variables collectively describe oxygen delivery (FiO2), pulmonary gas exchange efficiency (PaO2, SaO2), ventilatory adequacy (PaCO2, RespRate, MechVent), and the resulting systemic acid–base balance (pH). Grouping them lets the model learn the tightly coupled patterns that arise during hypoxemia, hypercapnia, mechanical ventilation adjustments, and respiratory failure—yielding a more coherent representation of a patient’s real-time respiratory status.
@@
    Singleton biomarkers
    Each remaining signal represents a distinct physiological domain (hepatic, renal, hematologic, hemodynamic, neurologic, metabolic, or cardiac). Their organ-specific pathophysiology favored treating them individually, preserving granular patterns while keeping the grouping scheme simple and interpretable.
8.2 CD
@@
This grouping reflects known biological relationships, enhances the interpretability of the model’s subset-level attributions, and improves performance compared to unstructured or purely univariate representations. It enables GMAN to exploit interactions among related features while maintaining a modular structure that aligns with clinical reasoning.
9 Experimental Setup, Hyperparameter Choices and Grouping Configurations
@@
9.1 General Experimental Setup
9.1.1 Crohn’s Disease Experiments
Unless otherwise noted, we trained all CD models for a maximum of 300 epochs using the Adam optimizer with weight decay in the 1e-2, 1e-5 range. We used a ReduceLROnPlateau scheduler with a max learning rate in the 1e-2, 1e-4 range, min learning rate in the 1e-7, 1e-8 range, factor in the 0.2-0.9 range and patience=100
We trained all GMAN models with batch size of range 64, 256, dropout in the 0.1-0.2 range, n_layers in the 2, 10 range, hidden_channels in the 4, 64 range, num_lab_ids_embed in the 5, 8 range, num_biom_embed in the 3, 5 range, num_units_embed in the 3, 5 range.
@@ 9.1.1 Crohn’s Disease Experiments
9.1.2 In-Hospital Mortality (P12) Experiments
We trained all PhysioNet12 models for a maximum of 500 epochs using the Adam optimizer with weight decay of 1e-4. We used a ReduceLROnPlateau scheduler with a max learning rate in the {1e-3, 1e-5} range, min learning rate of 1e-5, factor of 0.5 and patience=20.
We trained GMAN models with batch size of range {16, 32}, dropout of 0.2, n_layers in the {3, 5} range, hidden_channels in the {32, 64} range, num_lab_ids_embed in the {5, 8} range, num_biom_embed in the {3, 5} range, num_units_embed in the {3, 5} range.
9.1.3 Fake-News Detection (GosspiCop) Experiments
We trained all GNAM GosspiCop models for a maximum of 400 epochs using the Adam optimizer with weight decay of 1e-4. We used a ReduceLROnPlateau scheduler with a max learning rate in the {1e-3, 5e-5} range, min learning rate of 1e-8, factor of 0.5 and patience=20.
We trained GMAN models with batch size of 16, dropout in {0.0, 0.5} range, n_layers in the {3, 5} range, hidden_channels in the {16, 128} range.
Random seeds were fixed for reproducibility, and results are reported across three independent runs. All models were trained on a single NVIDIA Quadro RTX 8000 (48GB) GPU.
9.2 Grouping Configurations for Clinical Tasks
In both clinical tasks, the configuration of input feature subsets (i.e., how we grouped input biomarkers into multivariate trajectories) was treated as a manually tuned hyperparameter. These groupings determine how GMAN combines individual graph representations prior to final prediction, and they affect both the expressivity and interpretability of the model.
@@ 9.2 Grouping Configurations for Clinical Tasks
We compared GMAN’s performance under the following grouping strategies:
  • All singletons - 36 groups as the number of biomarkers.
@@
  • Flat grouping: All 17 biomarkers are modelled independently.
  • Biologically driven grouping (see Appendix 8.2): Biomarkers are grouped into 7 clinically coherent subsets (e.g., inflammation markers, immune cell subtypes, liver function).

### v3

@@ 5 Conclusion
We introduced Super Mixing Additive Networks (SuperMAN), a framework for learning from sets of graphs that represent irregular and asynchronous temporal signals. SuperMAN was designed to handle real-world scenarios where multiple signal types are collected at uneven intervals, such as medical records with heterogeneous blood tests or event logs in complex systems. The framework combined strong predictive performance with built-in interpretability, offering importance scores at the node, graph, and subset levels. SuperMAN allows practitioners to integrate domain priors when available, trading fine-grained interpretability for greater expressivity. Across experiments on real-world high-stakes tasks, SuperMAN achieved state-of-the-art results. Beyond predictive accuracy, SuperMAN’s interpretability capabilities proved particularly valuable in domains like healthcare, where uncovering phase transitions and providing actionable insights is critical.
Appendix A Appendix
Appendix B Theoretical Framework
B.1 Proof of Theorem 3.1
We will prove that SuperMAN is strictly more expressive than GNAN. To prove this, we use a ground truth function that is a feature-level XOR. Let a single-node graph be endowed with binary features $x=(x_{1},x_{2})\in\{0,1\}^{2}$ and define the target $f_{\oplus}(x)=x_{1}\oplus x_{2}$.
@@ Appendix B Theoretical Framework
Now we will show that SuperMAN can express $f_{\oplus}$. Place the two features in the same subset $F=\{x_{1},x_{2}\}$ and choose the subset-network
@@
For the four binary inputs this mapping returns $(0,1,1,0)$, exactly $f_{\oplus}$. Hence SuperMAN represents a function unattainable by GNAN, proving that SuperMAN is strictly more expressive.
B.2 Proof of Theorem 3.2
Let $S$ be a set of graphs $\{G_{i}\}_{j=1}^{m}$. Let $S_{1}=\{S_{i}\}_{i=1}^{m}$ be a partition of $S$ such that $|S_{i}|=1$. Let $S_{2}=\{S_{i}\}_{i=1}^{k}$ such that there exists $k$ with $|S_{k}|>1$. with a subset partition $\{S_{i}\}_{i=1}^{k}$. We will prove that a SuperMAN trained over $S_{2}$ is strictly more expressive than a SuperMAN trained over $S_{1}$.
@@
If each graph is placed in its own subset, SuperMAN aggregates additively: the model output is
@@
because the final SuperMAN stage simply sums subset scores :contentReference[oaicite:0]index=0:contentReference[oaicite:1]index=1. Write $a=\phi(0)$ and $b=\phi(1)$. To realise $f_{\oplus}$ via a threshold $\tau$ we would need
@@
Group the two graphs together and use a DeepSet $\Phi(S_{2})=g\!\bigl(\sum_{i=1}^{2}f(x_{i})\bigr)$ with $f(x)=x$ and g(s)=s(2s)g(s)=s(2-s). Then
@@
B.3 Four-Point Condition and Recoverability
We now turn to structural identifiability. We prove that when input graphs are connected, acyclic, and positively weighted (i.e., trees), the pairwise distance matrix learned by SuperMAN encodes the full structure of the graph, up to isomorphism. This provides theoretical justification for the model’s ability to reason over temporal structure without needing explicit graph supervision.
The following theorem shows that if the graph satisfies the four‑point condition (Buneman, 1974), SuperMAN can reconstruct the original graph from the transformed distance matrix that is fed to SuperMAN as the graph input:
Theorem B.1.
Let $G$ be a graph represented by an adjacency matrix $A$, and $D$ be the transformed distance-matrix for SuperMAN. Then if $D$ satisfies the four‑point condition, SuperMAN can learn $\rho$ such that $\rho(D)=A$.
@@ B.3 Four-Point Condition and Recoverability
Define the pair‑wise distance matrix $D\in\mathbb{R}^{n\times n}$ by
@@
where $P_{G}(u,v)$ is the unique uuvv path in $G$. Then:
@@
    Tree‑metric property. $D$ satisfies the four‑point condition of Buneman (Buneman, 1974); hence $(V,D)$ is a tree metric.
@@
Appendix C Computational Complexity and Efficient Implementation
In this section, we provide a big-O analysis of the time complexity of SuperMAN. We also provide an efficient approach to implement SuperMAN.
Computational Complexity
The computational complexity of SuperMAN is as follows:
  • Scales linearly with the number of graphs $m$: Each graph is processed independently or in small subsets, so total cost is $\mathcal{O}(m)$ assuming fixed per-graph cost.
  • Scales quadratically with the number of nodes per graph $n$: Due to the dense aggregation over all node pairs in ExtGNAN, the per-graph cost is $\mathcal{O}(n^{2})$.
  • Overall complexity for a set of graphs is:
$\mathcal{O}\!\left(m\cdot K\cdot n^{2}\cdot d_{\psi}\right)$
where $K$ is the number of feature groups and $d_{\psi}$ is the cost of evaluating the multivariate neural networks.
Adapting the complexity
We note that while the per-graph complexity of ExtGNAN is $O(n^{2})$ in the most general case, the distance function $\Delta$ can be masked to enforce any desired sparsity pattern between nodes, and therefore adapt to any complexity limitation, including a linear one. For example, one can define the $\Delta$ to only be the distance between adjacent nodes in a trajectory, and then it will be both linear in memory and time.
Efficient Implementations
In the main paper, we present SuperMAN with the objective of maximal clarity, e.g., by presenting vector entry-wise operations. Nonetheless, the operations of SuperMAN can be done in an optimized fashion for GPU, through tensor operations.
Appendix D Dataset Details
D.1 PhysioNet P12
@@
D.2 Crohn’s Disease Prediction
@@
D.2.1 Clinical Context and Related Work for predicting CD onset
@@
Several studies have assessed the predictive potential of standard clinical blood tests. For example, (Vestergaard et al., 2023) analyzed six routine biomarkers from 1,186 Danish patients eventually diagnosed with CD, achieving moderate predictive performance (AUROC of 0.74) approximately six months before clinical diagnosis. Larger-scale analyses, such as those using UK Biobank data, combined multiple standard biomarkers and basic demographic information, reporting similar predictive performances (AUROCs typically between 0.70–0.75). These analyses generally utilized methods like logistic regression, random forests, or gradient-boosted trees, favored for structured clinical datasets.
Other studies have integrated advanced biochemical data, known as multi-omics, including large-scale protein measurements (proteomics), metabolites (metabolomics), or genomic markers. (Garg et al., 2024) for instance, combined 67 blood biomarkers with approximately 2,900 plasma proteins from the UK Biobank, achieving an AUROC of 0.786. (Woerner et al., 2025) combined genetic risk scores with extensive proteomic data, achieving an AUROC of 0.76 for CD prediction up to five years prior to diagnosis. Similar multi-omics approaches employing microbiome profiling, immune signaling molecules (cytokines), or lipid molecules typically achieve AUROCs between 0.75 and 0.80 but often involve significant cost, specialized laboratory analyses, and reduced consistency across diverse patient cohorts.
@@
Appendix E Biomarker Subset Groupings
E.1 P12
@@
    Arterial blood gas profile
    [pH, PaCO2, PaO2, SaO2, HCO3, FiO2]
    This group captures systemic acid–base status (pH, HCO3), carbon dioxide clearance (PaCO2), oxygenation (PaO2, SaO2), and inspired oxygen fraction (FiO2). Together they form the canonical arterial blood gas panel, enabling the model to detect respiratory derangements such as hypoxemia, hypercapnia, or metabolic compensation.
@@
    Complete blood count
    [WBC, HCT, Platelets]
    This cluster summarizes hematologic composition by measuring leukocyte-mediated immune response (WBC), oxygen-carrying capacity (HCT), and clotting potential (Platelets). Joint modeling supports recognition of systemic inflammation, anemia, and coagulopathy.
  3. 3.
    Comprehensive metabolic panel
    [Glucose, Na, K, Mg, BUN, Creatinine]
    These biomarkers represent key substrates and electrolytes (Glucose, Na, K, Mg) and renal waste products (BUN, Creatinine). Grouping them provides a unified view of metabolic balance, electrolyte homeostasis, and kidney function.
  4. 4.
    Liver function tests
    [ALT, AST, ALP, Albumin, Bilirubin]
    These biomarkers assess hepatocellular injury (ALT, AST), cholestasis (ALP, Bilirubin), and hepatic synthetic function (Albumin). Their combined interpretation reflects multiple dimensions of liver health.
  5. 5.
    Lipid and cardiac markers
    [Cholesterol, TroponinI, TroponinT, HR]
    This group integrates lipid metabolism (Cholesterol), cardiac injury markers (Troponin I, Troponin T), and heart rate (HR). Together, they provide insight into cardiovascular stress, myocardial injury, and metabolic risk.
  6. 6.
    Blood pressure profiles
    [SysABP, DiasABP, MAP, NISysABP, NIDiasABP, NIMAP]
    These variables capture invasive and non-invasive arterial blood pressures, reflecting systemic hemodynamics. Grouping them enables the model to learn coherent pressure dynamics rather than treating each measurement in isolation.
  7. 7.
    Ventilation mechanics
    [RespRate, MechVent]
    This group reflects mechanical and physiological components of ventilation. Their joint dynamics provide context for interpreting respiratory compensation and ventilatory support.
  8. 8.
    Tissue perfusion
    [Lactate, Urine]
    Elevated lactate indicates anaerobic metabolism, while urine output tracks renal perfusion. Together they provide complementary signals of global tissue perfusion and shock severity.
  9. 9.
    Global status indicators
    [GCS, Temp]
    These variables capture overall neurologic responsiveness (GCS) and systemic temperature regulation (Temp), providing global context on patient stability and severity of illness.
E.2 CD
@@
This grouping reflects known biological relationships, enhances the interpretability of the model’s subset-level attributions, and improves performance compared to unstructured or purely univariate representations. It enables SuperMAN to exploit interactions among related features while maintaining a modular structure that aligns with clinical reasoning.
Appendix F Experimental Setup, Hyperparameter Choices and Grouping Configurations
@@
F.1 Hyper-Parameters
We trained all models with 100 epochs using the Adam optimizer with weight decay 1e-5. We used a ReduceLROnPlateau scheduler with a max learning rate in the 1e-2, 1e-4 range, min learning rate in the 1e-7, 1e-8 range, factor in the 0.2-0.9 range, and patience=100
We trained all models with batch size of range {16, 32}, dropout rate in {0.1, 0.2}, number of layers in the {3, 4, 5} range, hidden channels in the {32, 64} range.
@@ F.1 Hyper-Parameters
F.2 Grouping Configurations for Clinical Tasks
In both clinical tasks, the configuration of input feature subsets (i.e., how we grouped input biomarkers into multivariate trajectories) was treated as a manually tuned hyperparameter. These groupings determine how SuperMAN combines individual graph representations prior to final prediction, and they affect both the expressivity and interpretability of the model.
@@ F.2 Grouping Configurations for Clinical Tasks
We compared SuperMAN’s performance under the following grouping strategies:
  • No grouping: Each biomarker is in its own size one subset.
@@
  • Liver Panel - one group of the biomarkers: [ALT, AST, ALP, Albumin, Bilirubin, Cholesterol] and the rest are singletons.
  • Pathway-Based Grouping: Biomarkers are organised based on their molecular or mechanistic roles, grouping them by their function in metabolism, homeostasis, or cellular composition: Energy metabolism [Glucose, Cholesterol, Lactate]; nitrogen waste clearance [BUN, Creatinine, Urine]; protein synthesis and enzymes [Albumin, ALT, AST]; liver function and cholestasis [ALP, Bilirubin]; acid–base balance [pH, HCO3]; gas transport [PaO2, PaCO2, SaO2, FiO2]; mineral homeostasis [Na, K, Mg]; hematologic composition [HCT, Platelets, WBC]; cardiovascular dynamics [SysABP, DiasABP, MAP, NISysABP, NIDiasABP, NIMAP, HR]; respiratory mechanics [RespRate, MechVent]; and cardiac injury [TroponinI, TroponinT]. Global status indicators are grouped separately as [Temp, GCS].
  • Organ-System Grouping Biomarkers are organized around organ systems and clinical monitoring domains (respiratory support, cardiovascular dynamics etc: oxygenation support [SaO2, PaO2, FiO2, MechVent]; ventilation and acid–base balance [PaCO2, RespRate, pH, HCO3]; cardiovascular dynamics [HR, MAP, SysABP, DiasABP, NIMAP, NISysABP, NIDiasABP]; perfusion and renal function [Urine, Lactate, Creatinine, BUN]; hepatobiliary function [Bilirubin, ALT, AST, ALP, Albumin]; inflammation and coagulation [WBC, Platelets]; electrolyte and oxygen-carrying capacity [Na, K, Mg, HCT]; metabolic reserve [Glucose, Cholesterol]; myocardial injury [TroponinI, TroponinT]; and global status indicators [Temp, GCS].
@@
  • No grouping: Each biomarker is in its own size one subset.
  • Biologically driven grouping (see Appendix E.2): Biomarkers are grouped into 7 clinically coherent subsets (e.g., inflammation markers, immune cell subtypes, liver function).